# Frame In-N-Out: Unbounded Controllable Image-to-Video Generation

### v1

@@ 5.3 Unbounded Conditioning
Starting with the first frame condition $\mathit{I}_{0}$, we want the pixel-aligned motion intention to get over the border constraints. Hereby, we first extrapolate the first frame region to a larger area called the unbounded canvas region as shown in Fig. 3, which is defined by the provided top-left and bottom-right expansion pixels quantity $\mathbf{B}_{canvas}$. We define the expansion transformation from the original size of the first frame region to the canvas region coverage as $\tau_{canvas}(\cdot)$. The first frame is transformed $\tau_{canvas}(\mathit{I}_{0})$ by zero-padded. Then, we adjust the absolute and relative position encoding system by setting the top-left of the canvas region as the (0,0,0) index for temporal, horizontal, and vertical directions. Under this setting, our motion control signal can be expanded to any area inside the canvas region, which is also defined as $\mathbf{c}_{trajs}$.
@@
In the training, to accommodate faster training [64] without utilizing an attention mask for different resolutions between batches [11], we resize all videos to the same canvas resolution 384$\times$480 and the original first frame region is versatile based on the dataset curation from Sec. 4. Thanks to the flexible nature of the Diffusion Transformer with absolute and relative position embedding, we can set the canvas size arbitrarily in the inference with desirable results even if we train with a fixed size.
@@ 5.4 Unifying Identity Reference Conditions
Our model leverages the causal 3D VAE nature in modern video Diffusion Transformer [77], where not only sets of frames can be compressed, but one single image can be encoded. Therefore, we use the same pre-trained VAE $\mathcal{E}$ to encode the Identity (ID) reference $\mathbf{f}$. Inspired by Concat-ID [90] and recent progress in text-to-image generation with ID reference [51, 86], we resize and scale the ID reference to the same resolution as the canvas size $\mathbf{B}_{canvas}$ and then do frame-wise concatenation between the latent ID reference and the video frames. The ID reference will be added with random augment noise $n$ before VAE encode. To align the channel number, we add zero padding $\emptyset$ on the corresponding first frame and motion condition channels. The formula can be expressed as:
@@
where $\text{Concat}_{f}\left\{\cdot,\cdot,\cdots\right\}$ denotes frame-wise concatenation.
This method takes advantage of the 3D full attention nature of the video Diffusion Transformer. The text tokens, video tokens, and ID reference tokens will be token-wise concatenated after the patchification procedure and then jointly optimized together. Further, by reusing all well-trained normalization, projection, and feedforward modules, the training becomes more stable and implementation becomes more elegant. Though OmniControl [51] and EasyControl [86] might apply a shifted offset for the position encoding, this method does not provide a similar data distribution for the learned fixed position encoding on the ID reference part in the model like CogVideoX [77]. We empirically find that directly copying the position encoding of the first frame to the ID reference frame leads to better numerical results.
@@ 5.5 Training
Our training is composed of two stages. In the first stage, we include motion control based on the Eq. 2 with the text prompt captioned from [74] to learn the fundamental conditioning and adapt to the absolute position encoding modification we have done. Our loss in this stage can be formulated as:
$\mathcal{L}=\mathbb{E}_{\mathbf{z},\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t,\mathbf{c}}\left[||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{z}_{t},y,\mathit{I}_{0},\mathbf{c}_{trajs},t)||^{2}_{2}\right],$ (7)
where $t$ is the timestep, and $\mathbf{z}_{t}$ is the noisy latent at timestep $t$. During the inference, pure noise $\mathbf{z}_{T}$ is gradually denoised from timestep $T$ until timestep 00 to a clean latent $\mathbf{z}_{0}$. Then, it will be decoded by the pre-trained VAE decoder $\mathbb{D}$ to convert back to pixel space.
@@
$\mathcal{L}=\mathbb{E}_{\mathbf{z},\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t,\mathbf{c}}\left[||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\tau_{canvas}(\mathbf{z}_{t}),y,\tau_{canvas}(\mathit{I}_{0}),\mathbf{c}_{trajs},\mathbf{f},t)||^{2}_{2}\right],$ (8)
@@ 6 Experiment
6.1 Implementaion Details
The training dataset we consider includes OpenVid-1M [39], VidGen-1M [52], and subset of Webvid-10M [4]. The specific data and filtering statistics can be found in the appendix. We use the reserved subset of the OpenVid-1M dataset as our evaluation test set for Frame In-N-Out curation. Our training is on a total batch size of 8 for 36K iterations in each stage respectively. The training resolution, also the canvas resolution, is 384$\times$480 for two stages. All the video is curated, processed, and fetched in 12 FPS standards. We apply the learning rate warmup for each stage of training in the first 400 steps. The learning rate is 2e-5. We apply 8Bit Adam [12] to save the GPU memory. Our inference step is 50 with classifier-free guidance [20]. The first frame and text dropout ratio is 5% each to augment the classifier-free guidance in the inference. To make the motion pattern in the dataset stronger, we randomly doubled the duration fetched to simulate a speed-up with a probability of 20% and 33% respectively for each stage. We randomly drop ID reference with a probability of 15% in the stage2 training, where we only have Frame Out to consider.
@@ 6.2 Proposed Evaluation Dataset Overview
Since we are the first work focusing on the In-N-Out pattern in video generation, we define evaluation datasets and metrics as follows. Our evaluation is composed of two parts: Frame Out and Frame In with identity (ID) reference. Though we don’t require perfect Frame In and Frame Out patterns in our training scenarios, for the expression of a fair intention, we make the setting the hardest level in the evaluation test set. In this way, we curate 183 and 189 cases for perfect FrameIn and FrameOut as evaluation datasets. All Frame In evaluation datasets will come with one and only one ID reference image. The benchmark will be released for future study.
@@ 6.3 Evaluation Metrics
We evaluate the generative quality of video models using three widely adopted metrics: Fréchet Inception Distance [19] (FID), Fréchet Video Distance [55] (FVD), and Learned Perceptual Image Patch Similarity [83] (LPIPS). Beyond these generative metrics, we modify traditional tracking, segmentation, LLM evaluation, and identity-preserving metrics to fit the In-N-Out pattern.
Trajectory Error (Traj. Err.) evaluates the Euclidean distance of all trajectory points between the GT and the generated videos estimated by the Co-Tracker3 [27]. Different from trajectory error metrics proposed in [68, 70], our In-N-Out scenario considers the full canvas region for both GT and generated videos. Since the baseline method cannot generate pixels out of the first frame, they will be zero-padded to the same resolution as the GT, which refers to the canvas size. Lower scores indicate more aligned motion controllability. This metric is intended to leverage the low-level accuracy of the tracking when the object leaves and re-enters the scene.
@@
VSeg $\text{VSeg MAE}=\frac{1}{F\times H\times W}\sum_{i=1}^{F}\sum_{j=1}^{H}\sum_{j=k}^{W}\left|M_{gen}(i,j,k)-M_{gt}(i,j,k)\right|,$ (9)
where $F$, $H$, and $W$ refer to the frame number, video height, and width, and $M_{gen}$ and $M_{gt}$ refer to the segmentation area of the generated video and the Ground-Truth video estimated by SAM2 [43]. Generated videos without expansion capability will also be zero-padded. This metric is intended to calculate the accuracy of Frame In and Frame Out from a high-level semantic perspective.
Vision Language Model evaluation (VLM) utilizes a modern large vision language model, Qwen 2.5 VL 32B [74], with full video sequences and an instruction prompt to justify if the object gets out of the first frame or enter the first frame. We evaluate the ratio of correctness compared to the returns with GT video inputs. This metric is intended to align the overall subjective success rate analysis, and the higher the better.
Relative DINO (Rel. DINO) inherits the traditional DINOV2 [40] from VBench [21, 88] by calculating the cosine similarity between the ID reference and each video frame. However, we found that in our Frame In-N-Out setting, there exist multiple frames where the ID reference does not appear in the video frames at all, which leads to a very low similarity score. Thus, we first calculate the average DINO similarity score with each frame and then focus on the absolute relative difference to the Ground-Truth DINO results:
Relative $\text{Relative DINO}=\frac{\left|\frac{1}{F}\sum_{t=1}^{F}\langle d_{ID}\cdot d^{GEN}_{t}\rangle-\frac{1}{F}\sum_{t=1}^{F}\langle d_{ID}\cdot d^{GT}_{t}\rangle\right|}{\frac{1}{F}\sum_{t=1}^{F}\langle d_{ID}\cdot d^{GT}_{t}\rangle},$ (10)
[…]
@@
DragAnything [70] 48.73 607.44 0.462 41.24 0.0480 0.624
Image Conductor [33] 99.29 1154.86 0.528 42.72 0.0552 0.544
ToRA [87] 57.83 566.78 0.362 40.72 0.0750 0.603
@@ 6.4 Frame Out Comparisons
We consider SOTA motion controllable image-to-video (I2V) model, including DragAnything [70], Image Conductor [33], and ToRA [87]. For these baselines, their architecture does not support conditions of motion trajectory points outside the first frame; thus, we implement these points not appearing in the conditioning motion images. By default, all motions only apply one point and one object for a fair comparison. Further, we include our stage 1 motion-guided image-to-video generation results.
@@
Figure 4: Qualitative comparison on our benchmark dataset. In (a), we compare our model on Frame Out cases against DragAnything [70] and ToRA [87]. Both baselines fail to fully move the person outside the image boundaries, while our model successfully handles a complete exit. In (b), we evaluate Frame In scenarios against Phantom [36] and SkyReals-A2 [13]. Only our model can reach Frame In effect with the designated identity.
@@ 6.5 Frame In Comparisons
We define the task as a conditioning generation that requires an identity (ID) reference image with the first frame and motion trajectory. This task is not well-defined before this paper. Thus, there does not exist an appropriate strong baseline in the literature. Based on the conditioning we need, we found that the recent rising elements-to-video generation (E2V) is the closest fit, which can take conditions of ID reference and the first frame. While some other text-to-video generation works like Direct-A-Video [75] have motion control with ID reference, these works are not conditioned on the first frame, so we do not believe that they are a good fit to express unbounded controllable video generation concept we want to present in this paper. For the E2V, we consider Phantom [36] and SkyReels-A2 [13], which is based on Wan2.1 [58] Diffusion Transformer. Further, for a fair comparison, we re-generate a more motion descriptive text prompt by LLM [74] to compensate that E2V models cannot take in motion conditioning.
As shown in Tab. 3, our method significantly outperforms prior elements-to-video models across all key metrics, including FID, FVD, LPIPS, Traj. Err., VSeg MAE, and VLM accuracy, demonstrating superior visual quality and precise motion controllability in the Frame In setting. While our Rel. DINO score is slightly less (-0.23) than Phantom with motion prompts, this is primarily due to our model faithfully following motion guidance that occasionally moves the identity outside the canvas, affecting frame-wise similarity.
@@ 8 Acknowledgement
The authors gratefully acknowledge Research Computing at the University of Virginia for providing computational resources and technical support that made the results of this work possible. Zezhou Cheng acknowledges the Adobe Research Gift, as well as GPU support provided by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program and the National Artificial Intelligence Research Resource (NAIRR) Pilot program. The bird video frames used in Figure 1 and Figure 3 were adapted from the YouTube video “Videos for Cats to Watch” by Paul Dinning, and are used under fair use for academic and non-commercial purposes.
@@ Appendix B Dataset Curation Details
The detailed filtering statistics can be found in Tab. 4. For faster curation, we switch the order between the camera filter and panoptic segmentation. This is because camera detection by CUT3R [65] spends much more time than image-based panoptic segmentation by Oneformer [22]. Though WebVid [4] has an easy concept for each video, which makes it easier to curate, they have a watermark for each video and their resolution is low. Thus, we only consider around 1.5M videos of it to balance the dataset diversity and quality.
@@
In the Image Scoring, we randomly select 2 images from each video and then get the score of these two images by executing multiple image-based assessment models. We sort the score from the smallest to the highest on each metric and filter based on the human subjective perspective for different datasets. Thus, the filtering strength is different between datasets based on their characteristics. Further, we consider the overlap scenarios between different scoring elements. We consider image quality assessment by ClipIQA [63], with the range lowest 3%, 5%, 15% filtered for OpenVid [39], VidGen [52], and WebVid [4] respectively. We consider text detection from an open-source repository EasyOCR  [^2] https://github.com/JaidedAI/EasyOCR, which is based on [2]. We sort the area size of the detected text and then filter the highest 15%, 10%, and 5% respectively. We consider aesthetic assessment from NIMA [50] from the codebase of [8]. We filter the lowest 5%, 5%, and 10% respectively. For the image complexity assessment, we use IC9600 [14] to evaluate the complexity score where a lower value means less complexity, which lacks effective content to learn, and a higher value means more complexity, which has too many features to learn, like dense population or sightseeing. We filter both the lowest 10% and highest 5% for OpenVid, the lowest 5% and highest 10% for VidGen, and the lowest 5% and highest 10% for WebVid.
In the Scene Cut, we use TransNet V2 [47], which is based on the comparison results from Cosmos [1]. We filter any video that is detected with more than 1 scene. This ratio is not very high based on our observation. We think this is because the long video cases are already filtered.
In the Camera Filter, we use CUT3R [65], a state-of-the-art model for 3D reconstruction and camera pose estimation in dynamic video. However, evaluating 3D point positions jointly is computationally expensive, so we adopt the 224-resolution model and estimate camera poses over a 10-second window at 6 FPS. For each frame, we extract the predicted rotation, translation, and focal length. To sort videos based on camera motion intensity, we compute a score combining translational and rotational errors using the following formula:
@@
In the Panoptic Segmentation, we consider 22 objects of COCO dataset [35] detected by OneFormer [22] as the identity of interest, which includes person, bicycle, car, motorcycle, airplane, bus, train, truck, boat, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, sports ball, kite, and flower. We want the identity to be more diverse than regular human face-oriented identity-preserving-to-video [90, 80, 34] (IP2V), but less than elements-to-video domain [13, 36] that consider all genre, either motionable or static objects. The maximum duration of the I-Frame adjustment mentioned in the main paper is only 5% of the full length. If the I-Frame index is farther away from this, we use the original frame index counted. We discard objects that are too small, less than 4% of the area, and too big, which occupy more than 40% area of the image. The number of points sampled from K-means ranges from $[12,36]$ points based on the aforementioned area range. To avoid dense labeling with the same identity, we only allow at most 3 objects with the same label; otherwise, the video will be flittered.
In Video Captioning, there is no filtering. We apply QWen2.5-32B-VL-Instruct [74], which is the SOTA model for video captioning. We sampled only 1 frame per second with a resolution of 320$\times$448 to save computation resources. The instruction text prompt we use is: Please describe the video in 50 words. Only describe the temporal change of the video provided without describing the spatial information in the first frame provided. Only show the information with the highest confidence. Don’t use any words like gesture, gesturing. We apply captioning before Stage 1 motion-guided image-to-video generation training and the text prompt will be filtered with the rest of the procedure.
In Motion Filter, we first resize all videos to 384$\times$512, which is the resolution used on the CoTracker demo [27]. We sampled our video to 24 FPS but stored the result in 12 FPS, which is our training dataset fetched FPS. The start frame is the frame index from the panoptic segmentation section and the end frame is 49 frames from it, which is also our training duration for each selected video clip. If the cycle tracking errors at the first frame are larger than 4% of the number of pixels in height, this tracking point will be filtered. If more than 33% of points of an object are filtered in this way, the entire object is not considered.
In In-N-Out Filter, we consider various aspect ratio of 16:9, 3:2, 4:3, 5:4, 1:1, and 4:5 with probability of 35%, 30%, 20%, 13%, 1%, 1% respectively, where the minimum height is 60%, 60%, 65%, 65%, 75%, and 85% of its original height respectively. The top-left position is randomly generated, and the code base will check if the selected box can fit in the image resolution. If not, this bounding box will be filtered, and consider the next one. We will do this way 2000 times. If there is no ideal bounding box found, this video is filtered. All video clip that starts from the index of panoptic segmentation will be considered. The SAM2 [43] will be utilized after at least one ideal bounding box is found. We use SAM2 to further refine the tracking points outlier cases. Since SAM2 is also expensive, we did not deploy SAM2 with the CoTracker3 [27] at the motion filter stage. Instead, we only use it at the final stage for final improvement. We further discard identity reference images that are less than 10% of the image resolution size based on SAM2 estimation.
T
@@
Stage OpenVid [39] VidGen [52] WebVid [4]
@@
ConcatID [90]
Direct-A-Video [75]
ReVideo [38]
VideoAnyDoor [54]

### v2

@@ 5.3 Unbounded Conditioning
Starting with the first frame condition $\mathit{I}_{0}$, we want the pixel-aligned motion intention to get over the border constraints. Hereby, we first extrapolate the first frame region to a larger area called the unbounded canvas region as shown in Fig. 3, which is defined by the provided top-left and bottom-right expansion pixels quantity $\mathbf{B}_{canvas}$. We define the expansion transformation from the original size of the first frame region to the canvas region coverage as $\tau_{canvas}(\cdot)$. The first frame is transformed by zero-padded $\tau_{canvas}(\mathit{I}_{0})$. Then, we adjust the absolute and relative position encoding system by setting the top-left of the canvas region as the (0,0,0) index for temporal, horizontal, and vertical directions. Under this setting, our motion control signal can be expanded to any area inside the canvas region, which is also defined as $\mathbf{c}_{trajs}$.
@@
In the training, to accommodate faster training wang2025transpixar without utilizing an attention mask for different resolutions between batches dehghani2023patch , we resize all videos to the same canvas resolution, but the unmasked first frame region can be versatile based on the dataset curation strategy from Sec. 4. Thanks to the flexible nature of the Diffusion Transformer with absolute and relative position embedding, we can set the canvas size arbitrarily in inference with desirable results, even if we train with a fixed size.
@@ 5.4 Unifying Identity Reference Conditions
Our model leverages the causal 3D VAE nature in modern video Diffusion Transformer yang2024cogvideox , where not only sets of frames can be compressed, but one single image can be encoded. Therefore, we use the same pre-trained VAE $\mathcal{E}$ to encode the Identity (ID) reference $\mathbf{f}$. Inspired by Concat-ID zhong2025concat and recent progress in text-to-image generation with ID reference tan2024ominicontrol ; zhang2025easycontrol , we resize and scale the ID reference to the same resolution as the canvas size $\mathbf{B}_{canvas}$ and then do frame-wise concatenation between the latent ID reference and the video frames. The ID reference will be added with random augment noise $n$ before being encoded by VAE. To align the channel number, we add zero padding $\emptyset$ on the corresponding first frame and motion condition channels. The formula can be expressed as:
@@
where $\text{Concat}_{f}\left\{\cdot,\cdot,\cdots\right\}$ denotes frame-wise concatenation. After forwarding the Diffusion Transformer, text tokens and ID tokens will be discarded, without contributing to the loss.
This method leverages the 3D full attention nature of the video Diffusion Transformer. The text tokens, video tokens, and ID reference tokens will be token-wise concatenated after the patchification procedure and then jointly optimized together. Further, by reusing all well-trained normalization, projection, and feedforward modules, the training becomes more stable and the implementation becomes more elegant. Though OmniControl tan2024ominicontrol and EasyControl zhang2025easycontrol might apply a shifted offset for the position encoding, this method does not provide a similar data distribution for the learned fixed position encoding on the ID reference part in the model like CogVideoX yang2024cogvideox . We empirically find that directly copying the position encoding of the first frame to the ID reference frame leads to better numerical results.
@@ 5.5 Training
Our training is composed of two stages. In the first stage, we include motion control based on the Eq. 2 with the text prompt captioned from yang2024qwen2 to learn the fundamental conditioning and adapt to the absolute position encoding modification we have done. Our loss in this stage can be formulated as:
[…]
where $t$ is the timestep, and $\mathbf{z}_{t}$ is the noisy latent at timestep $t$. During the inference, pure noise $\mathbf{z}_{T}$ is gradually denoised from timestep $T$ until timestep 0 to a clean latent $\mathbf{z}_{0}$. Then, it will be decoded by the pre-trained VAE decoder $\mathbb{D}$ to convert back to pixel space.
@@
[…]
@@ 6 Experiment
6.1 Implementation Details
[…]
@@ 6.2 Proposed Evaluation Dataset Overview
Since we are the first work focusing on the In-N-Out pattern in video generation, we define evaluation datasets and metrics as follows. Our evaluation is composed of two parts: Frame Out and Frame In with identity (ID) reference. Though we don’t require perfect Frame In and Frame Out patterns in our training scenarios, for the expression of a fair intention, we set the setting to the hardest level in the curation of the evaluation test set. In this way, we curate 183 and 189 cases for ideal FrameIn and FrameOut as evaluation datasets. All Frame In evaluation datasets will come with one and only one ID reference image. The benchmark will be released for future study.
@@ 6.3 Evaluation Metrics
We evaluate the generative quality of video models using three widely adopted metrics: Fréchet Inception Distance heusel2017gans (FID), Fréchet Video Distance unterthiner2018towards (FVD), and Learned Perceptual Image Patch Similarity zhang2018unreasonable (LPIPS). Beyond these generative metrics, we modify traditional tracking, segmentation, LLM evaluation, and identity-preserving metrics to fit the In-N-Out pattern.
Trajectory Error (Traj. Err.) evaluates the Euclidean distance of all trajectory points between the GT and the generated videos estimated by the Co-Tracker3 karaev2024cotracker3 . Different from trajectory error metrics proposed in wang2024motionctrl ; wu2024draganything , our In-N-Out scenario considers the full canvas region for both GT and generated videos. Since the baseline method cannot generate pixels out of the first frame, they will be zero-padded to the same resolution as the GT, which refers to the canvas size. Lower scores indicate more aligned motion controllability. This metric is intended to leverage the low-level accuracy of the tracking when the object leaves and re-enters the scene.
@@
$\text{VSeg MAE}=\frac{1}{F\times H\times W}\sum_{i=1}^{F}\sum_{j=1}^{H}\sum_{j=k}^{W}\left|M_{gen}(i,j,k)-M_{gt}(i,j,k)\right|,$ (9)
where $F$, $H$, and $W$ refer to the frame number, video height, and width. $M_{gen}$ and $M_{gt}$ refer to the segmentation area of the generated video and the Ground-Truth video estimated by SAM2 ravi2024sam . The tracking points estimated by Co-Tracker3 served as the visual prompt for segmentation. Generated videos by the baseline methods without expansion capability will also be zero-padded. This metric is intended to calculate the accuracy of Frame In and Frame Out from a high-level semantic perspective.
Vision Language Model evaluation (VLM) utilizes a SOTA open-source vision language model, Qwen 2.5 VL 32B yang2024qwen2 to justify if there is any object gets out of the first frame or enters the first frame. The Frame In instruction prompt is Please check if the object enter the frame. Return a Yes/No as the only response. The Frame Out instruction prompt is Please check if the object leave the frame. Return a Yes/No as the only response. We evaluate the ratio of correctness compared to the returns with GT video inputs. If the reponse is not yes or no, we will skip that cases; however, we do not observe response different from these two designated outputs, which is thanks to strong capability of Qwen2.5. Due to computation concern, we evenly sample 14 frames from the full video sequence. The video is based on non-padded results, which does not consider the canvas region. This metric is intended to align the overall subjective success rate analysis, and the higher the better.
Relative DINO (Rel. DINO) inherits the traditional DINOV2 oquab2023dinov2 from VBench huang2024vbench ; zheng2025vbench by calculating the cosine similarity between the ID reference and each video frame for Frame In comparison. However, we found that in our Frame In-N-Out setting, there exists numerous frames that the ID reference does not appear in the video frames at all, which leads to a very low similarity score. Thus, we first calculate the average DINO similarity score between ID reference and each frame and then focus on the absolute relative difference to the Ground-Truth DINO results:
$\text{Relative DINO}=\frac{\left|\frac{1}{F}\sum_{t=1}^{F}\langle d_{ID}\cdot d^{GEN}_{t}\rangle-\frac{1}{F}\sum_{t=1}^{F}\langle d_{ID}\cdot d^{GT}_{t}\rangle\right|}{\frac{1}{F}\sum_{t=1}^{F}\langle d_{ID}\cdot d^{GT}_{t}\rangle},$ (10)
where $\langle\cdot\rangle$ is the dot product operation for calculating cosine similarity. The video is based on non-padded results, which does not consider the canvas region. The lower the score, the better.
@@
DragAnything wu2024draganything 48.73 607.44 0.462 41.24 0.0480 0.624
Image Conductor li2025image 99.29 1154.86 0.528 42.72 0.0552 0.544
ToRA zhang2024tora 57.83 566.78 0.362 40.72 0.0750 0.603
@@ 6.4 Frame Out Comparisons
We consider SOTA motion controllable image-to-video (I2V) model, including DragAnything wu2024draganything , Image Conductor li2025image , and ToRA zhang2024tora . For these baselines, their architecture does not support conditions of motion trajectory points outside the first frame; thus, we implement these points not appear in the conditioning motion images. By default, all motions only apply to one point and one object for a fair comparison. Further, we include our stage 1 motion-guided image-to-video generation results.
@@
Figure 4: Qualitative comparison on our benchmark dataset. In (a), we compare our model on Frame Out cases against DragAnything wu2024draganything and ToRA zhang2024tora . Both baselines fail to fully move the person outside the image boundaries, while our model successfully handles a complete exit. In (b), we evaluate Frame In scenarios against Phantom liu2025phantom and SkyReels-A2 fei2025skyreels . Only our model can reach Frame In effect with the designated identity.
@@ 6.5 Frame In Comparisons
We define the task as a conditioning generation that requires an identity (ID) reference image with the first frame and motion trajectory. This task is not well-defined before this paper. Thus, there does not exist an appropriate strong baseline in the literature. Based on the conditioning we need, we found that the recent rising elements-to-video generation (E2V) is the closest fit, which can take conditions of ID reference and the first frame. While some other text-to-video generation works like Direct-A-Video yang2024direct have motion control with ID reference, these works are not conditioned on the first frame, so we do not believe that they are a good fit to express the unbounded controllable video generation concept we want to present in this paper. For the E2V, we consider Phantom liu2025phantom and SkyReels-A2 fei2025skyreels , which is based on Wan2.1 wang2025wan Diffusion Transformer. Further, for a fair comparison, we re-generate a more motion descriptive text prompt by LLM yang2024qwen2 to compensate that E2V models cannot take in motion conditioning.
As shown in Tab. 3, our method significantly outperforms prior elements-to-video models across all key metrics, including FID, FVD, LPIPS, Traj. Err., VSeg MAE, and VLM accuracy, demonstrating superior visual quality and precise motion controllability in the Frame In setting. While our Rel. DINO score is slightly less (-0.23) than Phantom with motion prompts; this is primarily due to our model faithfully following motion guidance that occasionally moves the identity outside the canvas, affecting frame-wise similarity.
@@ 8 Acknowledgement
The authors acknowledge the Adobe Research Gift, the University of Virginia Research Computing and Data Analytics Center, Advanced Micro Devices AI and HPC Cluster Program, Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, and National Artificial Intelligence Research Resource (NAIRR) Pilot for computational resources, including the Anvil supercomputer (National Science Foundation award OAC 2005632) at Purdue University and the Delta and DeltaAI advanced computing resources (National Science Foundation award OAC 2005572). The bird video frames used in Figure 1 and Figure 3 were adapted from the YouTube video “Videos for Cats to Watch” by Paul Dinning, and are used under fair use for academic and non-commercial purposes.
@@ Appendix B Dataset Curation Details
The detailed filtering statistics can be found in Tab. 4. For faster curation, we switch the order between the camera filter and panoptic segmentation. This is because camera detection by CUT3R wang2025continuous spends much more time than image-based panoptic segmentation by Oneformer jain2023oneformer . Though WebVid bain2021frozen shows a clearer and more direct concept for each video than other datasets, they have watermarks, and their resolution is low. Thus, we only consider around 1.5M videos of it to balance the dataset diversity and quality.
@@
In the Image Scoring, we randomly select 2 images from each video and then get the score of these two images by executing multiple image-based assessment models. We sort the score from the smallest to the highest on each metric and filter based on the human subjective perspective for different datasets. Thus, the filtering strength is different between datasets based on their characteristics. Further, we consider the overlap scenarios between different scoring elements. We consider image quality assessment by ClipIQA wang2023exploring , with the range lowest 3%, 5%, 15% filtered for OpenVid nan2024openvid , VidGen tan2024vidgen , and WebVid bain2021frozen respectively. We consider text detection from an open-source repository EasyOCR  [^2] https://github.com/JaidedAI/EasyOCR, which is based on baek2019character . We sort the area size of the detected text and then filter the highest 15%, 10%, and 5%, respectively. We consider aesthetic assessment from NIMA talebi2018nima from the codebase of pyiqa . We filter the lowest 5%, 5%, and 10%, respectively. For the image complexity assessment, we use IC9600 feng2022ic9600 to evaluate the complexity score, where a lower value means less complexity, which lacks effective content to learn, and a higher value means more complexity, which has too many features to learn, like dense population or sightseeing. We filter both the lowest 10% and highest 5% for OpenVid, the lowest 5% and highest 10% for VidGen, and the lowest 5% and highest 10% for WebVid.
In the Scene Cut, we use TransNet V2 soucek2024transnet , which is based on the comparison results from Cosmos agarwal2025cosmos . We filter any video that is detected with more than 1 scene. This ratio is not very high based on our observation. We think this is because the long video cases are already filtered initially.
In the Camera Filter, we use CUT3R wang2025continuous , a state-of-the-art model for 3D reconstruction and camera pose estimation in dynamic video. However, evaluating 3D point positions jointly is computationally expensive, so we adopt the 224-resolution model and estimate camera poses over a 10-second window at 6 FPS. For each frame, we extract the predicted rotation, translation, and focal length. To sort videos based on camera motion intensity, we compute a score combining translational and rotational errors using the following formula:
@@
In the Panoptic Segmentation, we consider 22 objects of COCO dataset lin2014microsoft detected by OneFormer jain2023oneformer as the identity of interest, which includes person, bicycle, car, motorcycle, airplane, bus, train, truck, boat, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, sports ball, kite, and flower. We want the identity to be more diverse than regular human face-oriented identity-preserving-to-video zhong2025concat ; yuan2024identity ; li2024photomaker (IP2V), but less than elements-to-video domain fei2025skyreels ; liu2025phantom that consider all genre, either motionable or static objects. The maximum duration of the I-Frame adjustment mentioned in the main paper is only 5% of the full length. If the I-Frame index is farther away from this, we use the original frame index counted. We discard objects that are too small, less than 4% of the area, and too big, which occupy more than 40% area of the image. The number of points sampled from K-means ranges from $[12,36]$ points based on the aforementioned area range. To avoid dense labeling with the same identity, we only allow at most 3 objects with the same label; otherwise, the video will be flittered.
[…]
In In-N-Out Filter, we consider various aspect ratio of 16:9, 3:2, 4:3, 5:4, 1:1, and 4:5 with probability of 35%, 30%, 20%, 13%, 1%, 1% respectively, where the minimum height is 60%, 60%, 65%, 65%, 75%, and 85% of its original height respectively. The top-left position is randomly generated, and the code base will check if the selected box can fit in the image resolution. If not, this bounding box will be filtered, and consider the next one. We will do this way 2000 times. If there is no ideal bounding box found, this video is filtered. All video clip that starts from the index of panoptic segmentation will be considered. The SAM2 ravi2024sam will be utilized after at least one ideal bounding box is found. We use SAM2 to further refine the tracking points outlier cases. Since SAM2 is also expensive, we did not deploy SAM2 with the CoTracker3 karaev2024cotracker3 at the motion filter stage. Instead, we only use it at the final stage for final improvement. We further discard identity reference images that are less than 10% of the image resolution size based on SAM2 estimation.
@@
Stage OpenVid nan2024openvid VidGen tan2024vidgen WebVid bain2021frozen
@@
ConcatID zhong2025concat
Direct-A-Video yang2024direct
ReVideo mou2024revideo
VideoAnyDoor tu2025videoanydoor